# Single-photon scattering in giant-atom topological-waveguide-QED systems

### v1

@@ IV Conclusion
In conclusion, we have studied the single-photon scattering in an SSH waveguide coupled to either one or two two-level giant atoms. In both cases, we have derived the unified analytical expressions of the single-photon scattering amplitudes for different coupling configurations. We have found that the single-photon scattering is determined by the coupling configurations, coupling distances, atomic resonance frequency, and dimerization parameter. By analyzing the analytical expressions of the scattering amplitudes, we have found that, when all (any two or more) legs of the giant atoms are coupled to the sublattices of the same (different) type, the scattering behavior is independent of (depends on) the dimerization parameter.
For photons scattered by a single two-level giant atom, we have considered four different coupling configurations according to the coupling-point distributions. In particular, we have shown that the scattering spectra can show different line shapes by adjusting jointly the quantum interference effect and topological effect. In the case of two giant atoms, we have considered the single-photon scattering for sixteen coupling configurations. By comparing the Lamb shift, exchanging interaction, individual decay, and collective decay, we have further simplified these sixteen coupling configurations to six coupling cases. In particular, the single-photon scattering spectra can exhibit Lorentzian, super-Gaussian, EIT-like, and asymmetric Fano line shapes for some coupling configurations. The parameter ranges for the appearance of these line shapes have been discussed in detail, which can provide a valuable reference for the experimental study of giant-atom topological-waveguide-QED systems. These results show that the topological-waveguide-QED systems with giant atoms can exhibit a quantum switch for coherent single-photon transport. This work will pave the way for the study of controllable single-photon device in giant-atom topological-waveguide-QED systems.
@@ Acknowledgements.
J.-Q.L. was supported in part by National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, No. 11935006, and No. 12421005), National Key Research and Development Program of China (Grant No. 2024YFE0102400), and Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). X.-L.Y. was supported in part by Hunan Provincial Postgraduate Research and Innovation project (Grant No. CX20230463).

### v2

@@ IV Conclusion
In conclusion, we have studied the single-photon scattering in an SSH waveguide coupled to either one or two two-level giant atoms. In both cases, we have derived the unified analytical expressions of the single-photon scattering amplitudes for different coupling configurations. We have found that the single-photon scattering is determined by the coupling configurations, coupling distances, atomic resonance frequency, and dimerization parameter. By analyzing the analytical expressions of the scattering amplitudes, we have found that, when all (any two or more) legs of the giant atoms are coupled to the sublattices of the same (different) type, the scattering behavior is independent of (depends on) the sign of dimerization parameter.
For photons scattered by a single two-level giant atom, we have considered four different coupling configurations according to the coupling-point distributions. In particular, we have shown that the scattering spectra can show different line shapes by adjusting jointly the quantum interference effect and topological effect. In the case of two giant atoms, we have considered the single-photon scattering for sixteen coupling configurations. By comparing the Lamb shift, exchanging interaction, individual decay, and collective decay, we have further simplified these sixteen coupling configurations to six coupling cases. In particular, the single-photon scattering spectra can exhibit Lorentzian, super-Gaussian, EIT-like, and asymmetric Fano line shapes for some coupling configurations. The parameter ranges for the appearance of these line shapes have been discussed in detail, which can provide a valuable reference for the experimental study of giant-atom topological-waveguide-QED systems. These results show that the topological-waveguide-QED systems with giant atoms can not only exhibit a quantum switch for coherent single-photon transport, but also serve as a platform for probing the topological characteristics of these waveguide systems. This work will pave the way for the study of controllable single-photon devices in giant-atom topological-waveguide-QED systems.
@@ Acknowledgements.
J.-Q.L. was supported in part by the National Natural Science Foundation of China (Grants No. 12175061, No. 12247105, No. 11935006, and No. 12421005), the National Key Research and Development Program of China (Grant No. 2024YFE0102400), and the Hunan Provincial Major Sci-Tech Program (Grant No. 2023ZJ1010). X.-L.Y. was supported in part by the Hunan Provincial Postgraduate Research and Innovation project (Grant No. CX20230463).